[Brahms-l] Shift report 20050621 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 21 2005 - 00:24:53 EDT
Shifter leader was AG
Shifters were 


1. Around 1.5 hrs of beam.
Have had problems with supermon - it keept freezing just before I tried to print.
Checks with Supermon and TED during runs - showed no problems with detectors except for run 16702 which had no entries for TPM2 in row19 and 20.
In buliding 1002b the red light farthest to the left is blinking. There is a note near the light which says: Do not remove red plastic fence without written premission from the electrical system group head. George L. Ganetis: phone 4476/pager 4146. Have reported to JHL: No further action taken should be taken than to note it here since it isn't our equipment.
2. Fill Pattern: 111 X 111_p4, Polarization(10.00, 10.35 pm): B/Y: -45.07 +/- 2.66, -50.78 +/- 2.64:

Run 15700-15702

Trig2(212k), Trig3(63k), Trig4(37k) ,Trig5(262k)

Pedestal run 15696, Non pedestal run 15698

3. RD switched the drift voltage connection to the L1458A ps. from slot 7 ch 1 to slot 11 ch 4 TPM2 has drift voltage again.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 21 00:25:13 2005

This archive was generated by hypermail 2.1.8 : Tue Jun 21 2005 - 00:25:18 EDT